This equation has a { on the left side. I do not understand how to solve for x,y and z. I know I have to pair up equations but each attempt proves useless.
5x + 4y - z =1
2x - 2y + z = 1
- x - y + z = 2

Answer by scott8148(6628)

adding 1st eqn and 2nd eqn ___ 7x + 2y = 2 ___ multiplying by 3 ___ 21x + 6y = 6 (A)

adding 1st eqn and 3rd eqn ___ 4x + 3y = 3 ___ multiplying by -2 ___ -8x - 6y = -6 (B)

adding eqn (A) and eqn (B) ___ 13x = 0 ___ x = 0

substituting ___ 7(0) + 2y = 2 ___ y = 1

substituting ___ 5(0) + 4(1) - z = 1 ___ z = 3